# Functional Emotion Modeling in Biomimetic Reinforcement Learning

### v2

@@ Abstract
We explore a functionalist approach to emotion by employing an ansatz—an initial set of assumptions—that a hypothetical concept generation model incorporates unproven but biologically plausible traits. From these traits, we mathematically construct a theoretical reinforcement learning framework grounded in functionalist principles and examine how the resulting utility function aligns with emotional valence in biological systems. Our focus is on structuring the functionalist perspective through a conceptual network, particularly emphasizing the construction of the utility function, not to provide an exhaustive explanation of emotions. The primary emphasis is not of planning or action execution, but such factors are addressed when pertinent. Finally, we apply the framework to psychological phenomena such as humor, psychopathy, and advertising, demonstrating its breadth of explanatory power.
@@ I Introduction
An ansatz is a German term meaning ”approach” or ”starting point,” commonly used in physics, mathematics, and related fields. It refers to an educated guess or assumption about the form of a solution to a problem, particularly when solving complex equations or modeling physical systems. By making specific assumptions about the structure or behavior of a system, researchers can simplify intricate problems into more manageable forms. For example, in quantum mechanics, an ansatz might involve proposing a particular wavefunction form for a particle, while in statistical physics, it could involve assuming equilibrium conditions for a system. While an ansatz introduces simplifications, it is typically based on known physical principles or empirical observations to ensure its validity. This approach allows scientists and engineers to derive approximate solutions that capture the essential features of a problem without getting bogged down by overwhelming complexity.
@@ II-A Rationale of assumptions
Assumption (ii) is counterintuitive from an information density perspective. Given an n-bit bandwidth, more data can be represented if each permutation is allowed to represent a separate concept. However, this assumes that maintaining a 1 or 0 state require the same energy. In biological systems, this is simply not be true, hence the preference for sparseness [8]. From the analysis offered in Levy and Baxter [9], the cost of 10 spikes in 200 ms is more than 100 times that of maintaining the resting potential, and under those conditions the optimal fraction of active neurons is about 3%. Given this disparity, we can see how an encoding scheme satisfying assumption (ii) may increase evolutionary fitness, especially given the high energy cost of brain mass even with such optimizations.
@@ IV-C Extinction
The function of the S-UF is to identify and predict their associated P-UF components through learned experiences. However, due to the complexity of reality, these learned associations may not always be accurate. The question arises: how are incorrect associations eliminated and correct associations maintained? In contrast to P-UF, we define the S-UF such that depleted valence is not restored over time by default. If a positive S-UF component is repeatedly activated without concurrent activation of its associated P-UF component, the S-UF component will deplete entirely, leading to extinction. Conversely, if the activation of the P-UF component follows the activation of the S-UF component, the mapped valence of the S-UF component is restored through the same process that created it.
@@ IV-D Opposite Valence Interactions
Due to the relatively untargeted nature of unmodulated anger, it is also possible that the sources of negative valence are not specifically targeted initially. Rather, all active perceptrons representing active concepts outside of the threatened positive S-UF component are. Given that design, anger has a more global effect akin to that of P-UF activation. Targeted anger may be due to modulation from the DE module, which we will cover later.
@@ V-A Definitions
    $v_{i}(t)=v_{i}\cdot\text{active}(P_{i},t)$ (1)
    where $v_{i}$ is the intrinsic valence of $P_{i}$, and $\text{active}(P_{i},t)=1$ if active at time $t$, 0 otherwise.
@@
    S-UF Perceptrons: A set $\{s_{i}\mid i\in\mathbb{N}\}$, where a single perceptron is represented as $s_{i}$ (where $i$ indexes the perceptrons). Each has a binary activation state $f_{\text{char}}(i,t)=1$ if active at time $t$, 0 otherwise, and a utility value $UV_{i}(t)$ representing learned valence. A single S-UF component is composed of a subset of all perceptrons, $S_{i}\subseteq\{s_{j}\mid j\in\mathbb{N}\}$
@@
    $e_{i}(t)=\alpha\cdot e_{i}(t-1)+f_{\text{char}}(i,t)$ (2)
@@ V-C Collective Valence of S-UF Components
$UV_{S_{k}}(t)=\sum_{s_{i}\in S_{k}}UV_{i}(t)\cdot f_{\text{char}}(i,t)$ (4)
Here, only perceptrons active at time $t$ ($f_{\text{char}}(i,t)=1$) contribute to the component’s valence. This formulation captures partial activation: the valence experienced depends on how many perceptrons within $S_{k}$ are active, reflecting varying emotional intensity.
@@ V-D S-UF Valence Dynamics
$\text{diffusion}_{i}(t)=\beta\cdot\eta\cdot f_{\text{char}}(i,t)\sum_{j\sim i,f_{\text{char}}(j,t)}(UV_{j}(t)-UV_{i}(t))$ (6)
@@ V-G2 Anger Activation
Anger $A(t)$ activates when the loss of positive valence exceeds a threshold $\theta$:
@@
$A(t)=\text{clip}\left(\text{Loss}(t)-\theta,0,A_{\text{max}}\right)$ (13)
where $A_{\text{max}}$ is the maximum possible anger value.
Anger decays naturally over time. As the $\text{Baseline}(t)$ normalizes, $\text{Loss}(t)$ decreases, driving $A(t)$ lower.
@@ V-G3 Defensive Mechanisms
  • Reduced Valence Diffusion: The diffusion rate $\beta$ adjusts dynamically, limiting diffusion as anger increases:
    $\beta_{\text{effective}}(t)=\beta\cdot\left(1-\frac{A(t)}{A_{\text{max}}}\right)$ (14)
@@ V-G4 Limited Recovery Mechanism
$\text{max\_recovery}_{i}(t)=\text{max}((0,\text{cap}_{i}(t)$
$\text{EMS\_recovery}_{i}(t-1)))$ (17)
The desired recovery is:
@@ V-G5 Modulation of P-UF Valence Outputs
$v_{j}(t)=v_{j}^{\text{base}}(t)\cdot\left(1-k_{j}\cdot\frac{A(t)}{A_{\text{max}}}\right)$ (21)
@@ V-G7 Diffusion of Local Anger
$\text{diffusion}_{a_{i}}(t)=\beta\cdot\eta\cdot f_{\text{char}}(i,t)\cdot\sum_{j\sim i,f_{\text{char}}(j,t)}\left(a_{j}(t)-a_{i}(t)\right)$ (24)
@@ V-G8 Threat Neutralization
When the threat triggering anger is resolved (e.g. through external changes or internal reassessment) or enough time has passed, anger $A(t)$ naturally decreases as $\text{Loss}(t)$ diminishes due to baseline recalibration via the EMA. This reduction restores $\beta_{\text{effective}}(t)$ toward $\beta$, potentially allowing extinction of positive S-UF components if the threat persists without resolution. The DE module may further modulate anger- such as by adjusting global anger $A(t)$, adjusting local anger $a_{i}(t)$, or redirecting it to alternative targets. Under assumption (ii), valence can travel via diffusion along active perceptron chains that reflect conscious reasoning, such as linguistic reframing, based on the DE module’s situational assessment. However, the specific mechanisms of the DE-driven modulation are complex and context-dependent, warranting exploration in future work.
@@ VI-A7 Perceptron valence limit (Differing total valence capacity in concepts)
Suppose there are limitations on the valence that a single perceptron can hold, as described by the $L_{i}$ term in the formulations section. Consequently, concepts requiring more perceptrons to define would have a greater overall capacity for valence. If the complexity of a concept is generally proportional to the number of perceptrons needed to represent it, as implied by assumption (ii), then complex, multi-faceted concepts could accumulate higher levels of valence compared to simpler ones. This aligns with the intuition that simpler objects or entities, such as a toy dog, generally evokes less profound emotional responses than more complex entities, such as a real dog.

### v3

@@ Abstract
We explore a functionalist approach to emotion by employing an ansatz—an initial set of assumptions—that a hypothetical concept generation model incorporates unproven but biologically plausible traits. From these traits, we mathematically construct a theoretical reinforcement learning framework grounded in functionalist principles and examine how the resulting utility function aligns with emotional valence in biological systems. Our focus is on structuring the functionalist perspective through a conceptual network, particularly emphasizing the construction of the utility function, not to provide an exhaustive explanation of emotions. The primary emphasis is not on planning or action execution, but such factors are addressed when pertinent. Finally, we apply the framework to psychological phenomena such as humor, psychopathy, and advertising, demonstrating its breadth of explanatory power.
@@ I Introduction
An ansatz is a German term meaning ”approach” or ”starting point,” commonly used in physics, mathematics, and related fields. It refers to an educated guess or assumption about the form of a solution to a problem, particularly when solving complex equations or modeling physical systems. By making specific assumptions about the structure or behavior of a system, researchers can simplify intricate problems into more manageable forms. For example, in quantum mechanics, an ansatz might involve proposing a particular wavefunction form for a particle, while in statistical physics, it could involve assuming equilibrium conditions for a system. While an ansatz introduces simplifications, it is typically based on known physical principles or empirical observations to ensure its validity. This approach allows scientists and engineers to derive solutions that capture the essential features of a problem without getting bogged down by overwhelming complexity.
@@ II-A Rationale of assumptions
Assumption (ii) is counterintuitive from an information density perspective. Given an n-bit bandwidth, more data can be represented if each permutation is allowed to represent a separate concept. However, this assumes that maintaining a 1 or 0 state requires the same energy. In biological systems, this is simply not true, hence the preference for sparsity [21]. From the analysis offered in Levy and Baxter [22], the cost of 10 spikes in 200 ms is more than 100 times that of maintaining the resting potential, and under those conditions the optimal fraction of active neurons is about 3%. Given this disparity, we can see how an encoding scheme satisfying assumption (ii) may increase evolutionary fitness, especially given the high energy cost of brain mass even with such optimizations.
@@ IV-C Extinction
The function of the S-UF is to identify and predict its associated P-UF components through learned experiences. However, due to the complexity of reality, these learned associations may not always be accurate. The question arises: how are incorrect associations eliminated and correct associations maintained? In contrast to P-UF, we define the S-UF such that depleted valence is not restored over time by default. If a positive S-UF component is repeatedly activated without concurrent activation of its associated P-UF component, the S-UF component will deplete entirely, leading to extinction. Conversely, if the activation of the P-UF component follows the activation of the S-UF component, the mapped valence of the S-UF component is restored through the same process that created it.
@@ IV-D Opposite Valence Interactions
Due to the relatively untargeted nature of unmodulated anger, it is also possible that the sources of negative valence are not specifically targeted initially. Rather, all active perceptrons representing active concepts outside of the threatened positive S-UF component are. Given that design, initial anger has a more global effect akin to that of P-UF activation. Concepts can then be marked by the initial anger and diffusion, similar to S-UF construction. The DE module can modulate further through mechanisms such as appraisal and attention, resulting in a more targeted response.
@@ V-A Definitions
    $v_{i}(t)=v_{i}\cdot f_{p}(P_{i},t)$ (1)
    where $v_{i}$ is the intrinsic valence of $P_{i}$, and $f_{p}(P_{i},t)=1$ if active at time $t$, 0 otherwise.
@@
    S-UF Perceptrons: A set $\{s_{i}\mid i\in\mathbb{N}\}$, where a single perceptron is represented as $s_{i}$ (where $i$ indexes the perceptrons). Each has a binary activation state $f_{s}(i,t)=1$ if active at time $t$, 0 otherwise, and a utility value $UV_{i}(t)$ representing learned valence. A single S-UF component is composed of a subset of all perceptrons, $S_{i}\subseteq\{s_{j}\mid j\in\mathbb{N}\}$
@@
    $e_{i}(t)=\alpha\cdot e_{i}(t-1)+(1-\alpha)\cdot f_{s}(i,t)$ (2)
@@ V-C Collective Valence of S-UF Components
$UV_{S_{k}}(t)=\sum_{s_{i}\in S_{k}}UV_{i}(t)\cdot f_{s}(i,t)$ (4)
Here, only perceptrons active at time $t$ ($f_{s}(i,t)=1$) contribute to the component’s valence. This formulation captures partial activation: the valence experienced depends on how many perceptrons within $S_{k}$ are active, reflecting varying emotional intensity.
@@ V-D S-UF Valence Dynamics
$\text{diffusion}_{i}(t)=\beta\cdot\eta\cdot f_{s}(i,t)\sum_{j\sim i}(UV_{j}(t)-UV_{i}(t))\cdot f_{s}(j,t)$ (6)
@@ V-G2 Anger Activation
Anger $A(t)=A_{p}(t)+A_{s}(t)$, where $A_{p}(t)$ is the global/primary anger term and $A_{s}(t)$ is the local/secondary anger term. $A_{p}(t)$ activates when the loss of positive valence exceeds a threshold $\theta$:
@@
$A_{p}(t)=\text{max}\left(\text{Loss}(t)-\theta,0\right)$ (13)
$A_{s}(t)$ is the sum of local anger scalars $a_{i}(t)$ among all active perceptrons:
$A_{s}(t)=\sum_{i\in\mathbb{N}}a_{i}(t)\cdot f_{s}(i,t)$ (14)
Primary anger decays naturally over time. As the $\text{Baseline}(t)$ normalizes, $\text{Loss}(t)$ decreases, driving $A_{p}(t)$ lower.
@@ V-G3 Defensive Mechanisms
  • Reduced Valence Diffusion: The diffusion rate $\beta$ as related to normal UV flow adjusts dynamically, limiting diffusion as anger increases:
    $\beta_{\text{effective}}(t)=\beta\cdot\left(1-\frac{\text{min}(A(t),A_{\text{thres}})}{A_{\text{thres}}}\right)$ (15)
    where $A_{\text{thres}}>0$ is the threshold of total anger beyond which P-UF inputs and normal UV diffusion inputs are frozen.
@@ V-G4 Limited Recovery Mechanism
$\text{max\_recovery}_{i}(t)=\text{clip}(\sigma+\text{max\_recovery}_{i}(t-1)$
$\text{EMS\_recovery}_{i}(t-1),0,\text{cap}_{i}(t))$ (18)
where $\sigma\geq 0$ is the rate at which the fallback recovery capacity is restored. The desired recovery is:
@@ V-G5 Modulation of P-UF Valence Outputs
$v_{j}(t)=v_{j}^{\text{base}}(t)\cdot\left(1-k_{j}\cdot\frac{\text{min}(A(t),A_{\text{thres}})}{A_{\text{thres}}}\right)$ (22)
@@ V-G7 Diffusion of Local Anger
$\text{diffusion}_{a_{i}}(t)=\beta\cdot\eta\cdot f_{s}(i,t)\cdot\sum_{j\sim i}\left(a_{j}(t)-a_{i}(t)\right)\cdot f_{s}(j,t)$ (25)
@@ V-G8 Threat Neutralization
When the threat triggering anger is resolved (e.g. through external changes or internal reassessment) or enough time has passed, anger $A(t)$ naturally decreases as $\text{Loss}(t)$ diminishes due to baseline recalibration via the EMA. This reduction restores $\beta_{\text{effective}}(t)$ toward $\beta$, potentially allowing extinction of positive S-UF components if the threat persists without resolution. The DE module may further modulate anger- such as by adjusting global anger $A(t)$, adjusting local anger $a_{i}(t)$, or redirecting it to alternative targets. Under assumption (ii), valence can travel via diffusion along active perceptron chains that reflect conscious reasoning, such as linguistic reframing, based on the DE module’s situational assessment. However, the specific mechanisms of the DE-driven modulation such as appraisal or attention are complex and context-dependent, warranting exploration in future work.
@@
V-H Toy Simulation
To illustrate the core valence dynamics (diffusion, depletion, anger protection), we implemented a minimal 4-concept toy world (food, home, intruder, pain) in Python. The code is available in a public repository [36], and the output is shown in Fig. 1. Food and pain are both tied to positive and negative P-UF components, respectively.
Figure 1: Example output of the 4-concept simulation showing valence trajectories under safe eating, intruder attack, and recovery phases.
@@ VI-A7 Perceptron valence limit (Differing total valence capacity in concepts)
Suppose there are limitations on the valence that a single perceptron can hold, as described by the $L_{i}$ term in the formulations section. Consequently, concepts requiring more perceptrons to define would have a greater overall capacity for valence. If the complexity of a concept is generally proportional to the number of perceptrons needed to represent it, as implied by assumption (ii), then complex, multi-faceted concepts could accumulate higher levels of valence compared to simpler ones. This aligns with the intuition that simpler objects or entities, such as a toy dog, generally evoke less profound emotional responses than more complex entities, such as a real dog.